OpenStudy (anonymous):

An initial investment of $2,100 earns 7% interest compounded continuously. What will the investment be worth in 16 years? (Round your answer to the nearest cent.)

OpenStudy (anonymous):

Pe^rt = 2100e^(.07)(16)

OpenStudy (anonymous):

$ 6199.55

OpenStudy (anonymous):

I got that part, I just got the calculation wrong. And @sheg, It was not correct

OpenStudy (anonymous):

I did the same thing you did, it was still wrong.

OpenStudy (anonymous):

well, equation that I wrote was just to compute the interest. To compute the value of the investment at maturity, you would need to add the interest calculation to the principle.

OpenStudy (anonymous):

can you show me how to get my answer?

OpenStudy (anonymous):

p(1+e^rt)

OpenStudy (anonymous):

Pe^kt

OpenStudy (anonymous):

how come it is not correct ??????????????

OpenStudy (anonymous):

I don't know, Webassign just said it was wrong

OpenStudy (anonymous):

just cross check it is correct according to ur question

OpenStudy (anonymous):

@indy25, What is the real answer? I need a good answer.

OpenStudy (anonymous):

I don't know what the answer is.... I've got it wrong

OpenStudy (anonymous):

\[A = P \times ( 1+ r)^n = $ 2100 ( 1 + 0.07)^{16 } = $ 2100 (2.952163749) = $ 6199.543872\]

OpenStudy (anonymous):

now u guys tell me where it is wrong !!

OpenStudy (anonymous):

it's compounded continously, your series is incorrect.

OpenStudy (anonymous):

it's an exponential series

OpenStudy (anonymous):

I am getting 6436.19+2100=8536.19=pe^rt

OpenStudy (anonymous):

pe^rt +p = p(1+e^rt)

OpenStudy (anonymous):

both mean the same @ indy

OpenStudy (anonymous):

P(1+r)^n = p/1-(1+r) if 0<n<infinity

OpenStudy (anonymous):

prove it.